OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x + 3)(x – 1)(x + 5)? 3, 1, –5 –3, 1, –5 3, –1, 5 –3, –1, 5

OpenStudy (anonymous):

if any one of those 3 sets of parentheses has a zero value, then the whole function is zero. So what x values create a value of zero inside any one of the sets of parentheses? Like the first one... (x+3)... what "x" makes (x+3) = 0 ?

OpenStudy (anonymous):

wut?

OpenStudy (anonymous):

the first part of the function is (x+3)... if you evaluate f(x) at x = -3, then that first part becomes (-3 + 3) = 0 and so it doesn't matter what the rest of the parts of the function are, f(-3) is going to be 0.

OpenStudy (anonymous):

so does that leave only a and d?

OpenStudy (anonymous):

x = -3 appears in b and d, not a and d... Yes, so then for the other two "zeros", just look at the "x" value you need to make the 2nd and 3rd parenthesis equal 0. (x-1) = 0 <------- x = ? (x+5) = 0 <-------x = ?? Those x values are the other two "zeros" of the function.

OpenStudy (anonymous):

yea sorry, so D!

OpenStudy (anonymous):

no, "b" :) x must be 1 to make (x-1) = 0, and x must be -5 to make (x + 5) = 0 It's just backwards from the way you were thinking...

OpenStudy (anonymous):

@andreadesirepen

OpenStudy (anonymous):

oh, im dumb. thanks!!

OpenStudy (anonymous):

not dumb, just in a hurry, I bet... glad to help :)